OpenStudy (anonymous):

(x+2)/(x−4)≤ 0 Solve and graph.

OpenStudy (anonymous):

1/(x^2+2x-4x-8)

OpenStudy (anonymous):

1/(x^2-2x-8)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

X+2>=0 and x-4<0 x>=-2 and x<4 solution [-2,4) OR x+2<=0 and x-4>0 x<=-2 and x>4 solution is empty set

OpenStudy (anonymous):



OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

nevermind. its not.

OpenStudy (anonymous):

you always can check solution by pluging it into inequality